# The infrared luminosity of retired and post-starburst galaxies: A cautionary tale for star formation rate measurements

### v1

@@ 4 Results
The lower left panel colours the galaxies by the fraction of total dust luminosity that is fit by the ambient ISM dust component in magphys ($f_{\mu,\mathrm{IR}}$), as opposed to the birth cloud dust surrounding the star-forming regions. Here we see a more distinct discontinuity between the blue and red sequences, with the majority of red sequence galaxies requiring $\lesssim 10$% additional contribution from a hotter birth cloud component, while the blue sequence galaxies have typically $\gtrsim 50$% contribution from hotter birth clouds.
The lower right panel colours the galaxies by the temperature of the ambient ISM dust component fit by magphys. Constraining dust temperature accurately requires more dust bands than a simple dust luminosity, we therefore required galaxies to be detected at $5\sigma$ in 3 or more 22–500 µm bands for this analysis (2,877 galaxies). While there is a significant scatter, the predominant trend is for galaxies with higher $L_{\mathrm{H}\,\alpha}$ to have higher temperature ISM dust ($\theta_{p}=-15^{\circ}$). There is a small residual trend with stellar mass, but not to the same extent as for $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$. This shows that the changing ISM dust temperature is not the cause of the changing luminosity ratio. magphys also fits for the temperature of the birth cloud thermal component (not shown). We find a similar result, with galaxies with higher $L_{\mathrm{H}\,\alpha}$ having higher temperature birth cloud dust and any residual trend with stellar mass being small.
@@ 5 Discussion
However, $L_{\mathrm{TIR}}$ and $L_{\mathrm{H}\,\alpha}$ respond differently to the stellar population mix. Because $L_{\mathrm{H}\,\alpha}$ emission caused by HOLMES is very weak, $L_{\mathrm{H}\,\alpha}$ drops very sharply once the star-forming regions vanish. On the other hand, $L_{\mathrm{TIR}}$ remains high because dust can be heated by all stars emitting in the UV and optical wavelength regime. Thus, the $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio increases, reflecting the higher fraction of dust luminosity coming from ambient ISM compared to the fraction of $L_{\mathrm{H}\,\alpha}$ coming from older stars. We hypothesise that it is this that causes the smooth trend between $L_{\mathrm{TIR}}$, $L_{\mathrm{H}\,\alpha}$ and $M_{*}$ continuing into the red sequence, without discontinuity, despite the apparently different ionising and heating mechanisms, and star formation histories of the galaxies. The changing $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ with sSFR then simply reflects the differing relative importance of OB stars to HOLMES for $L_{\mathrm{H}\,\alpha}$, and diffuse ISM to birth clouds for $L_{\mathrm{TIR}}$ emission.
As discussed in Section 3, when luminosities are converted into a SFR, a constant SFR over a fixed timescale is assumed (10 Myr for $L_{\mathrm{H}\,\alpha}$ and 100 Myr for $L_{\mathrm{TIR}}$). While this might be a reasonable approximation for a star-forming galaxy, it clearly is not for a young post-starburst galaxy where the SFR has dropped precipitously over the past hundred Myr. To untangle the relative importance of star formation history versus assumed star-formation timescale to the excess $L_{\mathrm{TIR}}$ seen in retired, post-starburst and star-forming galaxies with $W_{\mathrm{H}\,\alpha}<10$, we turn to toy models.
@@ 5.1 Toy SED models
To illustrate how $L_{\mathrm{H}\,\alpha}$ and $L_{\mathrm{TIR}}$ depend on a galaxy’s stellar populations, we constructed a set of toy SED models based on average star formation histories for star-forming, post-starburst and retired galaxies in the starlight–SDSS main galaxy sample catalogue. We note that the post-starburst galaxies were selected via the Goto (2007) method, meaning they have completely ceased star-formation and thus match the subset of our post-starburst galaxies with $W_{\mathrm{H}\,\alpha}<3$. The models were constructed in mass bins of 1 dex using Bruzual & Charlot (2003) models with a Chabrier (2003) initial mass function. Further details are provided in Appendix A.
@@
where $\tau_{V}$ is the effective dust attenuation in the $V$-band and $q_{\lambda}\equiv\tau_{\lambda}/\tau_{V}$ is the shape of the dust attenuation law, here taken to be the Cardelli et al. (1989) extinction curve with $R_{V}=3.1$. We created models with four values of the effective dust attenuation, $\tau_{V}=0.1,0.5,1.0$ and $2.0$. $L_{\mathrm{TIR}}$ was then calculated by integrating the difference between the intrinsic and dust reddened spectrum, and assuming all light attenuated by dust is re-radiated in the mid- and far-IR [^10] Changing the attenuation law to a Charlot & Fall (2000) two-component dust law using $q_{\lambda}=(\lambda/{5500\,\mathrm{\AA}})^{-0.7}$, and for a birth cloud $V$-band attenuation twice that of the ISM, results in a modest increase in the predicted $\log L_{\mathrm{TIR}}$ of 0.16 dex for the SF and of 0.08 dex for the post-starburst and retired galaxies..
@@
When going from models for star-forming galaxies to post-starburst/retired galaxies, within the same mass and $\tau_{V}$ bin (fixed symbol style and size, changing colour) there is a slight decrease in $L_{\mathrm{TIR}}$ as old stars still heat the diffuse ISM, and a huge drop in $L_{\mathrm{H}\,\alpha}$. This highlights the fact that HOLMES are a much less powerful ionising source than OB stars: the rate of ionising photons decreases by five orders of magnitude from a 10 Myr to a 1–10 Gyr stellar population; see figure 1, panel b of Cid Fernandes et al. (2011).
@@
The slightly higher $L_{\mathrm{TIR}}$ of the quenched post-starburst galaxies compared to retired galaxies is due to the presence of a large number of relatively hot A and F stars to heat the dust, rather than the recent precipitous drop in SFR (which from Figure 4 can be seen to be insignificant for well over the 100 Myr averaging time assumed for $L_{\mathrm{TIR}}$ in star-forming galaxies). The comparatively small difference between post-starburst and retired galaxies, compared to the difference between post-starburst and star-forming galaxies, shows that the precise stellar population balance is a much less important factor than the current specific star formation rate in determining the $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio.
@@
Our toy models do not include post-starburst galaxies that are in the process of quenching, which are seen in Figure 1 to lie just above the star-forming sequence. Here we expect a combination of a rapidly declining recent SFR and a reduction in birth cloud dust to be contributing to the enhanced $L_{\mathrm{TIR}}$ compared to $L_{\mathrm{H}\,\alpha}$ (see also Hayward et al., 2014).
@@
5.2 Dust-star geometry, interstellar radiation field and cloud density
@@
Interestingly, the many complications that could arise in terms of dust temperature, dust-star geometry, dust cloud density and even ionisation source do not appear to be important to our results, at least to first order, even for post-starburst galaxies which should provide a test of the most extreme quenching conditions. At far-IR wavelengths, the temperature of the dust is thought to be determined mostly by the dust-star geometry, rather than the shape of the interstellar radiation field (ISRF, e.g. Bernard et al., 2010). Thus, for any given stellar population mix, the stars which emit the most UV–optical photons will contribute most to the total $L_{\mathrm{TIR}}$, i.e. OB stars in star-forming galaxies, AF stars in post-starburst galaxies, or G and later in retired galaxies, but the differing temperature of the stars is not relevant. Our results in Figure 2 suggest that dust temperature predominantly depends on $L_{\mathrm{H}\,\alpha}$, with a small residual mass trend. Galaxies with the highest $L_{\mathrm{H}\,\alpha}$ in our sample have higher ambient ISM and dense BC dust temperatures, and this is expected to be due to the closer proximity of the dust to the star-forming regions, rather than the changing ISRF (e.g. Nersesian et al., 2019).
@@
The lack of any appreciable difference in the temperature of post-starburst galaxy ISM dust compared to normal galaxies is interesting (bottom right panel of Figure 2), as we might expect the A/F stars to be co-located with the majority of the ISM dust following the end of the starburst, and therefore able to enhance the heating in a similar way to the brightest starbursts. Rowlands et al. (2015) found a decrease in post-starburst dust temperature with time since burst. However, they were studying generally younger post-starburst galaxies than in this sample, and temperature was estimated from colour index rather than the magphys fit as used here. It seems likely that the dust heating observed in the youngest post-starbursts in their sample was due to a changing balance of hot birth cloud dust to cooler ISM dust following the star-formation episode, rather than a more general heating of the ambient ISM i.e. closer to the $f_{\mu,\mathrm{IR}}$ parameter in the bottom left panel of Figure 2.
@@ 5.3 SFR estimators and full spectral fitting
Another argument is that it does not matter where the photons come from, they still represent star formation, and should be counted even if the star formation happened long ago. This is because combining $L_{\mathrm{TIR}}$ with a UV luminosity to account for the unextincted photons accounts for most photons generated by stars that are still alive at the time of observation, regardless of their mass or when they were formed. However, this argument is only plausibly useful when integrating in both time and space, not when trying to estimate the ongoing SFR of a particular galaxy, or set of galaxies as it does not tell us the number of stars recently formed.
@@
For $L_{\mathrm{H}\,\alpha}$ the time-averaging is the main sequence lifetime of O and B stars, about 10 Myr. It is this short timescale that makes it attractive as a robust SFR indicator. For FIR, the time averaging is around 100 Myr for star-forming galaxies, which accounts for lower mass stars also heating the dust (Kennicutt, 1998). All conversions between a proxy luminosity and SFR require the assumption of a constant SFR over the averaging timeframe. For some galaxies, such as those which are currently rapidly declining in SFR, the longer estimators are therefore less suitable. Our results, supplemented by our toy models, highlight the subtle difference between star formation history versus contamination by ambient ISM dust in enhancing $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$. The post-starburst galaxies in our sample that are currently decreasing in SFR, but still have appreciable SFR, have a slightly enhanced $L_{\mathrm{TIR}}$ for their $L_{\mathrm{H}\,\alpha}$, and they lie close to ordinary star-forming galaxies. This could be due to their rapidly declining star formation history, as well as their reducing contribution of birth cloud dust to $L_{\mathrm{TIR}}$. On the other hand, the post-starburst galaxies that have completely quenched their SFR have a very substantially enhanced $L_{\mathrm{TIR}}$ for their $L_{\mathrm{H}\,\alpha}$, causing them to lie close to the retired galaxies. When the observed luminosity is dominated by photons from significantly lower mass stellar populations, we are no longer measuring the number of recently formed stars, but rather the integrated mass formed during the entire lifetime of the galaxy. This leads to a strong correlation between $L_{\mathrm{TIR}}$ and $M_{*}$, rather than $L_{\mathrm{TIR}}$ and SFR. This suggests that we should see different partial correlation angles between $L_{\mathrm{TIR}}$, $M_{*}$ and $L_{\mathrm{H}\,\alpha}$ for star-forming, retired and post-starburst galaxies. Unfortunately the small numbers of retired and post-starburst galaxies in our samples made it impossible to tease these relations apart, and a larger set of retired and quenched post-starburst galaxies would be required.
@@ 5.4 Comparison to previous work
Although papers studying the SFR calibrators commonly investigate galaxy types, few provide measurements of the $W_{\mathrm{H}\,\alpha}$ or sSFR of the galaxies being used for calibration. One exception is Hao et al. (2011) where the majority of galaxies have $W_{\mathrm{H}\,\alpha}>10$, and their figure 7 shows an overestimation of the SFR using an FIR-based calibrator by $0.3$ dex for the small number of galaxies with $W_{\mathrm{H}\,\alpha}>5$. This is entirely consistent with our results.
Rosario et al. (2016) looked at $L_{\mathrm{H}\,\alpha}$ versus $L_{\mathrm{TIR}}$ in an early Herschel-SDSS sample similar to that studied here, and while they did not link their results to star formation histories, their plots show the same excess of $L_{\mathrm{TIR}}$ in low sSFR galaxies. It is reassuring that an entirely different analysis leads to the same quantitative results.
Our results are also consistent with those of Nersesian et al. (2019) using the SED modelling code CIGALE (Boquien et al., 2019), who noted that galaxies with $\log_{10}\mathrm{(sSFR/yr}^{-1})>-10.5$ had dust mainly heated by radiation emitted by the young population, whereas lower sSFR had dust heated by the older population. This corresponds to $W_{\mathrm{H}\,\alpha}>10$, matching our results.

### v2

@@ 4 Results
The lower left panel colours the galaxies by the fraction of total dust luminosity that is fit by the ambient ISM dust component in magphys ($f_{\mu,\mathrm{IR}}$), as opposed to the birth cloud dust surrounding the star-forming regions. We note that this result is dependent on the magphys model, unlike previous results, but is useful to further interpret our results. Here we see a more distinct discontinuity between the blue and red sequences, with the majority of red sequence galaxies requiring $\lesssim 10$% additional contribution from a hotter birth cloud component, while the blue sequence galaxies have typically $\gtrsim 50$% contribution from hotter birth clouds.
The lower right panel colours the galaxies by the temperature of the ambient ISM dust component fit by magphys. Constraining dust temperature accurately requires more dust bands than a simple dust luminosity, we therefore required galaxies to be detected at $5\sigma$ in 3 or more 22–500 µm bands for this analysis (2,877 galaxies) [^10] For both the bottom left and bottom right panels, we verified that the results do not change if we require significant detections in more mid and far-IR bands.. While there is a significant scatter, the predominant trend is for galaxies with higher $L_{\mathrm{H}\,\alpha}$ to have higher temperature ISM dust ($\theta_{p}=-15^{\circ}$). There is a small residual trend with stellar mass, but not to the same extent as for $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$. This shows that the changing ISM dust temperature is not the cause of the changing luminosity ratio. magphys also fits for the temperature of the birth cloud thermal component (not shown). We find a similar result, with galaxies with higher $L_{\mathrm{H}\,\alpha}$ having higher temperature birth cloud dust and any residual trend with stellar mass being small.
@@ 5 Discussion
There are two scenarios by which $L_{\mathrm{TIR}}$ and $L_{\mathrm{H}\,\alpha}$ should not track each other, which are interrelated, yet subtly different. Firstly, dust heating of older stars contributes more to $L_{\mathrm{TIR}}$ than to $L_{\mathrm{H}\,\alpha}$. The extent to which this matters depends on (a) the mass of older stars in the galaxy, and therefore on the star formation history of the galaxy in the distant past, and (b) the presence of ambient ISM that is available to be heated. Secondly, even in a star-forming galaxy, if the SFR rapidly decreases, $L_{\mathrm{H}\,\alpha}$ will decrease rapidly while $L_{\mathrm{TIR}}$ will remain high for $\sim$100 Myr. In this section we argue that dust heating by older stars, rather than a recent change in SFR, is the dominant effect.
5.1 A two-component ISM
@@
However, $L_{\mathrm{TIR}}$ and $L_{\mathrm{H}\,\alpha}$ respond differently to the stellar population mix. Because $L_{\mathrm{H}\,\alpha}$ emission caused by HOLMES is very weak, $L_{\mathrm{H}\,\alpha}$ drops very sharply once the star-forming regions vanishes. On the other hand, $L_{\mathrm{TIR}}$ drops more slowly over a time period of 100 Myr, and remains persistently higher than $L_{\mathrm{H}\,\alpha}$ until the dust is destroyed, because dust can be heated by all stars emitting in the UV and optical wavelength regime (see figure 3 of Hayward et al. 2014). Thus, the $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio increases and remains high after a decline in star formation, reflecting the higher fraction of dust luminosity coming from the ambient ISM compared to the fraction of $L_{\mathrm{H}\,\alpha}$ coming from older stars. We hypothesise that it is this effect that causes the smooth trend between $L_{\mathrm{TIR}}$, $L_{\mathrm{H}\,\alpha}$ and $M_{*}$ continuing into the red sequence, without discontinuity, despite the apparently different ionising and heating mechanisms, and star formation histories of the galaxies. The changing $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ with $L_{\mathrm{H}\,\alpha}$/$M_{*}$ then simply reflects the differing relative importance of OB stars to HOLMES for $L_{\mathrm{H}\,\alpha}$, and diffuse ISM to birth clouds for $L_{\mathrm{TIR}}$ emission.
As discussed in Section 3, when luminosities are converted into a SFR, a constant SFR over a fixed timescale is assumed (10 Myr for $L_{\mathrm{H}\,\alpha}$ and 100 Myr for $L_{\mathrm{TIR}}$). While this might be a reasonable approximation for a star-forming galaxy, it clearly is not for a young post-starburst galaxy where the SFR has dropped precipitously over the past hundred Myr. To untangle the relative importance of star formation history versus ambient dust heating to the excess $L_{\mathrm{TIR}}$ seen in retired, post-starburst and star-forming galaxies with $W_{\mathrm{H}\,\alpha}<10$, we turn to toy models.
@@ 5.2 Toy SED models
To illustrate how $L_{\mathrm{H}\,\alpha}$ and $L_{\mathrm{TIR}}$ depend on a galaxy’s stellar populations, we constructed a set of toy SED models based on average star formation histories for star-forming, post-starburst and retired galaxies in the starlight–SDSS main galaxy sample catalogue. We note that the post-starburst galaxies were selected via the Goto (2007) method, meaning they have completely ceased star-formation and thus match the subset of our post-starburst galaxies with $W_{\mathrm{H}\,\alpha}<3$. The models were constructed in mass bins of 1 dex using Bruzual & Charlot (2003) models with a Chabrier (2003) initial mass function. Further details are provided in Appendix A, including a figure demonstrating the difference in star formation histories between the star-forming, retired and post-starburst samples (Figure 5).
@@
where $\tau_{V}$ is the effective dust attenuation in the $V$-band and $q_{\lambda}\equiv\tau_{\lambda}/\tau_{V}$ is the shape of the dust attenuation law, here taken to be the Cardelli et al. (1989) extinction curve with $R_{V}=3.1$. We created models with four values of the effective dust attenuation, $\tau_{V}=0.1,0.5,1.0$ and $2.0$. $L_{\mathrm{TIR}}$ was then calculated by integrating the difference between the intrinsic and dust reddened spectrum, and assuming all light attenuated by dust is re-radiated in the mid- and far-IR.
@@
When going from models for star-forming galaxies to post-starburst/retired galaxies, within the same mass and $\tau_{V}$ bin (fixed symbol style and size, changing colour) there is a slight decrease in $L_{\mathrm{TIR}}$ as old stars still heat the diffuse ISM, and a huge drop in $L_{\mathrm{H}\,\alpha}$. This highlights the fact that HOLMES are a much less powerful ionising source than OB stars: the rate of ionising photons decreases by five orders of magnitude from a 10 Myr to a 1–10 Gyr stellar population (see figure 1, panel b of Cid Fernandes et al. 2011).
@@
The slightly higher $L_{\mathrm{TIR}}$ of the quenched post-starburst galaxies compared to retired galaxies is due to the presence of a large number of relatively hot A and F stars to heat the dust, rather than the recent precipitous drop in SFR (which from Figure 5 can be seen to be insignificant for well over the 100 Myr averaging time assumed for $L_{\mathrm{TIR}}$ in star-forming galaxies). The comparatively small difference between post-starburst and retired galaxies, compared to the difference between post-starburst and star-forming galaxies, shows that the recent star formation history is a much less important factor than the specific star formation rate in determining the $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$ ratio.
@@
Our toy models do not include post-starburst galaxies that are in the process of quenching, which are seen in Figure 2 to lie just above the star-forming sequence. Here we expect a combination of a rapidly declining recent SFR and a reduction in birth cloud dust to be contributing to the enhanced $L_{\mathrm{TIR}}$ compared to $L_{\mathrm{H}\,\alpha}$ (see also Hayward et al., 2014). We also do not include any timescale for dust destruction, so our toy models are only pertinent to galaxies which have some ambient ISM available to be heated.
@@
5.3 Dust-star geometry, interstellar radiation field, dust destruction and cloud density
@@
Interestingly, the many complications that could arise in terms of dust temperature, dust-star geometry, dust destruction and dust cloud density and even ionisation source do not appear to be important to our results, at least to first order, even for post-starburst galaxies which should provide a test of the most extreme quenching conditions. At far-IR wavelengths, the temperature of the dust is thought to be determined mostly by the dust-star geometry, rather than the shape of the interstellar radiation field (ISRF, e.g. Bernard et al., 2010). Thus, for any given stellar population mix, the stars which emit the most UV–optical photons will contribute most to the total $L_{\mathrm{TIR}}$, i.e. OB stars in star-forming galaxies, AF stars in post-starburst galaxies, or G and later in retired galaxies, but the differing temperature of the stars is not relevant. The magphys SED modelling results shown in Figure 3 suggest that dust temperature correlates most strongly with $L_{\mathrm{H}\,\alpha}$, with a small residual mass trend. magphys modelling also suggests that galaxies with the highest $L_{\mathrm{H}\,\alpha}$ in our sample have higher ambient ISM and dense birth cloud dust temperatures. This is expected to be due to the closer proximity of the dust to the star-forming regions, rather than the changing ISRF (e.g. Nersesian et al., 2019). These particular results will likely be sensitive to the assumed model, and deserve further exploration.
@@
The lack of any appreciable difference in the temperature of post-starburst galaxy ISM dust compared to normal galaxies is interesting (bottom right panel of Figure 3), as we might expect the A/F stars to be co-located with the majority of the ISM dust following the end of the starburst, and therefore able to enhance the heating in a similar way to the brightest starbursts. Rowlands et al. (2015) found a decrease in post-starburst dust temperature with time since burst. However, they were studying generally younger post-starburst galaxies than in this sample, and temperature was estimated from colour index rather than the magphys fit as used here. It seems likely that the dust heating observed in the youngest post-starbursts in their sample was due to a changing balance of hot birth cloud dust to cooler ISM dust following the star-formation episode, rather than a more general heating of the ambient ISM i.e. closer to the $f_{\mu,\mathrm{IR}}$ parameter in the bottom left panel of Figure 3. Over time and with no subsequent star formation we expect the dust to be destroyed, and here post-starburst galaxies provide a valuable timescale that can be used to constrain the mechanisms responsible for dust destruction. Li et al. (2019) used the unique clock provided by post-starburst galaxies to measure a timescale for dust destruction of $\sim$250 Myr. By 1 Gyr post-burst, the majority of post-starburst galaxies have no detectable dust.
@@ 5.4 SFR estimators and full spectral fitting
Another argument is that it does not matter where the photons come from, they still represent star formation, and should be counted even if the star formation happened long ago. This is because combining $L_{\mathrm{TIR}}$ with a UV luminosity to account for the unextincted photons accounts for most photons generated by stars that are still alive at the time of observation, regardless of their mass or when they were formed. However, this argument is only plausibly useful when integrating in both time and space, not when trying to estimate the ongoing SFR of a particular galaxy, or set of galaxies, as it does not tell us the number of stars recently formed.
@@
For $L_{\mathrm{H}\,\alpha}$ the time-averaging is the main sequence lifetime of O and B stars, about 10 Myr. It is this short timescale that makes it attractive as a robust SFR indicator. For FIR, the time averaging is around 100 Myr for star-forming galaxies, which accounts for lower mass stars also heating the dust (Kennicutt, 1998). All conversions between a proxy luminosity and SFR require the assumption of a constant SFR over the averaging timeframe. For some galaxies, such as those which are currently rapidly declining in SFR, the longer estimators may be less suitable. Our results, supplemented by our toy models, highlight the subtle difference between star formation history versus contamination by ambient ISM dust in enhancing $L_{\mathrm{TIR}}$/$L_{\mathrm{H}\,\alpha}$. The post-starburst galaxies in our sample that are currently decreasing in SFR, but still have appreciable SFR, have a slightly enhanced $L_{\mathrm{TIR}}$ for their $L_{\mathrm{H}\,\alpha}$, and they lie close to ordinary star-forming galaxies. This could be due to their rapidly declining star formation history, as well as their reducing contribution of birth cloud dust to $L_{\mathrm{TIR}}$. On the other hand, the post-starburst galaxies that have completely quenched their SFR have a very substantially enhanced $L_{\mathrm{TIR}}$ for their $L_{\mathrm{H}\,\alpha}$, causing them to lie close to the retired galaxies. When the observed luminosity is dominated by photons from significantly lower mass stellar populations, we are no longer measuring the number of recently formed stars, but rather the integrated mass formed during the entire lifetime of the galaxy. This leads to a strong correlation between $L_{\mathrm{TIR}}$ and $M_{*}$, rather than $L_{\mathrm{TIR}}$ and SFR. This suggests that we should see different partial correlation angles between $L_{\mathrm{TIR}}$, $M_{*}$ and $L_{\mathrm{H}\,\alpha}$ for star-forming, retired and post-starburst galaxies. Unfortunately the small numbers of retired and post-starburst galaxies in our samples made it impossible to tease these relations apart, and a larger set of retired and quenched post-starburst galaxies would be required.
@@ 5.5 Comparison to previous work
It has long been known that dust heating by old stars will impact low sSFR galaxies more than their high sSFR counterparts, with many cautionary notes in the literature (see e.g. Kennicutt, 1998). In this section we compare our results to previous studies where mid- and far-IR luminosities have been noticed to apparently over-estimate SFR in galaxies.
Although papers purposefully studying SFR calibrators commonly investigate galaxy types, few provide measurements of the $W_{\mathrm{H}\,\alpha}$ or sSFR of the galaxies being used for calibration. One exception is Hao et al. (2011) where the majority of galaxies have $W_{\mathrm{H}\,\alpha}>10$, and their figure 7 shows an overestimation of the SFR using an FIR-based calibrator by $0.3$ dex for the small number of galaxies with $W_{\mathrm{H}\,\alpha}>5$. This is entirely consistent with our results.
Salim et al. (2016) presented the GALEX-SDSS-WISE Legacy catalogue, noting that galaxies with $\log_{10}\mathrm{(sSFR/yr}^{-1})<-11$ have mid-IR SFRs biased high by up to 2 dex, attributing this to the dust being heated primarily by old stars. Similarly, Rosario et al. (2016) looked at $L_{\mathrm{H}\,\alpha}$ versus $L_{\mathrm{TIR}}$ in an early Herschel-SDSS sample similar to that studied here, and while they did not link their results to star formation histories, their plots show the same excess of $L_{\mathrm{TIR}}$ in low sSFR galaxies. Kouroumpatzakis et al. (2023) used the SED modelling code CIGALE (Boquien et al., 2019) to show that mid-IR-based SFR estimators overpredict SFR compared to CIGALE for galaxies with $W_{\mathrm{H}\,\alpha}$ $<10$Å. It is reassuring that entirely different analyses have lead to the same quantitative results.
Our results are also consistent with those of Nersesian et al. (2019) who used CIGALE to note that galaxies with $\log_{10}\mathrm{(sSFR/yr}^{-1})>-10.5$ had dust mainly heated by radiation emitted by the young population, whereas lower sSFR galaxies had dust heated by the older population. This corresponds to $W_{\mathrm{H}\,\alpha}>10$, matching our results. Recently Parente et al. (2024) have used a semi-analytic galaxy evolution model coupled to the GRASIL radiative transfer model to demonstrate the dust content of galaxies transitioning through the green-valley remains high following star formation quenching, leading to relatively large 250 µm fluxes in agreement with GAMA observations.